OpenStudy (anonymous):

Find the value of g when 5g + 9 equals 2g + 15. B. explain how you can check that the value you found for g is correct. If your check does not work, does that mean that your result is incorrect? Explain.

OpenStudy (haseeb96):

so first 5 g +9 = 2g +15 okay ? as given in the question statement now 5g-2g =15 -9 3g =6 g =6/3 g=2 okay now put g=2 in the above equation(5g +9 = 2g +15) 5(2) +9 = 2(2) +15 10 +9 =4 +15 19 = 19 so it is showing that value is same on both sides so it is correct

OpenStudy (haseeb96):

do u have any problem ?

OpenStudy (haseeb96):

the value of g =2 and for checking i have done that how to check either your answer is correct or incorrect

OpenStudy (anonymous):

Yea I have more. Ty

OpenStudy (haseeb96):

okay bring them i will shake them lol